### v2

Abstract
Foundation models like CLIP and SAM have advanced computer vision and medical imaging via low-shot transfer learning, aiding CADD with limited data. However, their deployment faces two key challenges. distribution shift where pre-training and post-training data distributions differ (e.g., due to inter-center image acquisition) and confidence misalignment, which leads to overconfident errors. These issues surface differently, vision-language models (e.g., CLIP) suffer from 2D embedding shift (image-text misalignment), while medical models (e.g., SAM) encounter 3D domain shifts (e.g., scanner variation) and voxel-wise calibration need. Existing solutions are domain-specific. We propose StaRFM, a fusion of Fisher information penalty (FIP) and confidence misalignment penalty (CMP) tackling both challenges. It applies FIP, extended to 3D via patch-wise regularization, to reduce embedding shift, and CMP, reformulated for voxel-level predictions, to calibrate segmentation uncertainty. We derive PAC-Bayes bounds. FIP controls generalization via the Fisher-Rao norm, and CMP reduces calibration error via Brier score minimization. StaRFM surpasses baselines by +3.5% accuracy and 28% lower ECE on 19 vision datasets (e.g., ImageNet, Office-Home), achieves +4.2% DSC over SAM-FT and 4.8mm HD95 on medical benchmarks (e.g., BraTS, ATLAS), and reduces cross-domain gaps by up to 20%. The framework is plug-and-play, requiring minimal architectural changes. Code and models are available at: StaRFM.
keywords:
[…]
1 Introduction
Recent advances in foundation models such as CLIP Radford et al. (2021), Align Abdul Samadh et al. (2024) and Flamingo Alayrac et al. (2022), in vision-language and Segment Anything Model Kirillov et al. (2023), MedSAM Ma et al. (2024), BioMedCLIP Zhang et al. (2023), UNET Ronneberger et al. (2015), Dinov2 Oquab et al. (2023), and CXR-CLIP You et al. (2023) in medical image analysis have revolutionized machine learning by enabling zero-shot transfer learning to downstream tasks. However, their real-world deployment is hindered by two pervasive intertwined challenges. Distribution shift where test data diverge from training distribution $P_{tr}(y,x)\neq P_{ts}(y,x)$, and confidence misalignment, where models produce overconfident yet incorrect predictions Guo et al. (2017); Pereyra et al. (2017). These issues are particularly acute in safety-critical domains, from autonomous systems to medical diagnostics, where unreliable uncertainty estimates can lead to catastrophic failures Yao et al. (2024); Mirza et al. (2021).
In vision-language models (VLMs) like CLIP, Covariate shift where the source and target feature distributions differ ($P_{tr}(x)\neq P_{ts}(x)$) Shimodaira (2000); Sugiyama et al. (2007) while $P(y|x)$ remains unchanged. This shift is ubiquitous in real-world settings such as cross-domain visual classification, population-biased medical imaging, or protocol variations across clinical institutions Khan et al. (2025). Concurrently, such shifts exacerbate confidence misalignment, where softmax probabilities fail to reflect true likelihoods Murugesan et al. (2025).
Recent work Khan et al. (2025) proposed CalShift, a unified framework addressing both challenges via Fisher information regularization and a confidence misalignment penalty (CMP). While effective for 2D classification, its applicability to high-dimensional, spatially structured data (e.g., 3D medical images) remains unexplored. In this work, we extend CalShift Khan et al. (2025) published in CVPR2025, from 2D classification to vision-language and 3D medical imaging tasks.
Similar challenges emerge in medical foundation models like SAM Kirillov et al. (2023), when adapted to brain MRI segmentation, SAM’s predictions grow overconfident under domain shifts induced by scanner protocols or population biases Huang et al. (2024). This mirrors the covariate-confidence duality in VLMs but introduces unique complexities: volumetric data continuity, partial volume effects, and ambiguous anatomical boundaries Isensee et al. (2021). Current solutions either ignore calibration Chen et al. (2024) or rely on post-hoc fixes Wang (2023), leaving the core issue of joint robustness and reliability unresolved during training.
While both VLMs and medical foundation models face similar challenges of distribution shift and confidence misalignment, the solutions developed for each domain have remained siloed. In VLMs, techniques like Fisher regularization Khan et al. (2025) have proven effective for 2D image-text alignment but fail to account for the spatial continuity and anatomical constraints inherent to 3D medical data. Conversely, medical imaging approaches Huang et al. (2024) often focus on architectural adaptations without leveraging the theoretical foundations of distribution shift correction.
(a) Misaligned predictions caused by covariate shift.
(b) Aligned predictions after integration of CMP and FIP $I(\theta)$.
Figure 1: Workflow of the proposed StaRFM framework: The sub-figure (a) illustrates the confidence misalignment problem caused by covariate shift. The (bottom right) part of the sub-figure (a) show misaligned predictions. The subfigure (b) middle section represents the two components as recipe in method: Fisher information penalty ($I(\theta)$) for covariate shift correction and confidence misalignment penalty (CMP) for calibration. Both FIP and CMP are integrated into the CLIP text encoder. The final loss, $\mathcal{L}_{\text{StaRFM}}=\mathcal{L}_{\text{CLIP}}+\lambda_{1}I(\theta)+\lambda_{2}\text{CMP}$, combines the original CLIP loss with FIP and CMP to produce robust and aligned predictions, as shown in (bottom right) as output.
Contributions
In this work, we bridge this gap by generalizing the CalShift framework to 3D medical segmentation and by offering a unified perspective on domain shift and calibration across vision-language and medical domains which we refer as StaRFM. Our contributions are:
  1. 1.
    We extend CalShift’s Fisher information penalty (FIP) to $3$D data by computing patch-wise gradients over volumetric embeddings to add robustness to MRI domain shifts. We reformulate the CMP component at voxel level, penalizing overconfident predictions in lesion segmentation.
  2. 2.
    We derive new generalization bounds for segmentation tasks , showing how FIP tightens the PAC-Bayes gap under covariate shift, while CMP controls calibration error via Brier score minimization.
  3. 3.
    Through extensive experiments on 19 vision datasets , 8 domain shift and 3 medical benchmarks (BraTS, ATLAS), we demonstrate that our method consistently improves performance metrics such as accuracy, DSC, and HD95, achieving up to a +3.5% increase in accuracy and a 28% reduction in Expected Calibration Error (ECE) compared to the baselines.
  4. 4.
    StaRFM operates as a plug-and-play module for the text encoder used with foundation models such as CLIP and SAM, requiring minimal architectural changes and training effort.
By unifying robustness and calibration under a common framework, our work facilitates the development of trustworthy AI systems in high-stakes environments. In vision-language modeling, it resolves the long-standing issue of overconfidence under distribution shift in few-shot adaptation Ovadia et al. (2019). To enhance clarity, we delineate the working mechanism of our method. Figure 1 presents the overall workflow for vision adaptation tasks, encompassing both calibration and generalization under covariate shift.
In medical imaging, it presents the first training-time strategy for calibrated 3D segmentation under domain shift Jungo and Reyes (2019). Figure 2 represents the workflow of our method for medical segmentation tasks.
Figure 2: Workflow of the proposed StaRFM framework for medical segmentation tasks. Both FIP and CMP are integrated into the SAM prompt encoder. The final loss, $\mathcal{L}_{\text{StaRFM}}=\mathcal{L}_{\text{SAM}}+\lambda_{1}I_{3D}(\theta)+\lambda_{2}\text{CMP}_{3D}$, combines the original SAM loss with FIP and CMP to produce robust and aligned predictions.
2 Methods in comparison
Foundation models in vision and medicine. Foundation models like CLIP Radford et al. (2021), Flamingo Alayrac et al. (2022), SAM Kirillov et al. (2023), and MedSAM Ma et al. (2024) have transformed computer vision and clinical applications via zero-shot transfer. CLIP’s contrastive image-text pretraining enables strong downstream classification Zhou et al. (2022), while SAM’s prompt-based segmentation generalizes well on natural images. However, medical adaptations (e.g., MedSAM Ma et al. (2024), BioMedCLIP Zhang et al. (2023)) face challenges under distribution shifts. Such as, covariate shift arises when CLIP’s pretrained features $P_{tr}(x)$ diverge from target distributions $P_{ts}(x)$ Khan et al. (2024), while in medical imaging SAM exhibits overconfident errors on MRI and CT scans due to domain gaps with natural images Huang et al. (2024).
Covariate shift in vision-language and medical tasks. Covariate shift appears differently in vision-language models (VLMs) and medical image segmentation. In CLIP-like architectures, it occurs when pretrained text-image feature distributions $P_{src}(x)$ diverge from downstream task data $P_{tgt}(x)$ Khan et al. (2024). This divergence often results from dataset fragmentation in few-shot learning Zhou et al. (2022) and domain gaps between web-scale pretraining (e.g., LAION) and specific target domains Abdul Samadh et al. (2024). While prior work quantifies this shift using Fisher information $I(\theta)$ Khan and Syed (2025), it overlooks the complexities of 3D data.
In medical segmentation tasks (e.g., SAM on MRI), shift arises due to scanner and resolution variations across institutions Menze et al. (2014), as well as mismatched intensity distributions between natural and medical images Huang et al. (2024). Existing domain adaptation methods Chen et al. (2024) employ adversarial training to mitigate this, but often sacrifice model calibration. We address both issues by introducing regularization term which are integrated into loss function given in eq. 16.
Domain adaptation under covariate shift. Distribution shift is addressed differently across domains. In vision-language models, Fisher Information regularization Khan et al. (2024) mitigates CLIP’s covariate shift without requiring target labels, though it overlooks calibration. Adversarial approaches Wang et al. (2024), by contrast, involve extensive retraining.
In medical imaging, test-time adaptation Chen et al. (2024) tackles scanner-induced variations, while feature alignment Kamnitsas et al. (2017) typically relies on labeled target data. Our framework unifies these strategies by extending Fisher information to 3D medical patches $I_{3D}(\theta)=-\mathbb{E}_{x\sim P_{MRI}}\left[\frac{\partial^{2}\log P(x;\theta)}{\partial\theta^{2}}\right]$. This formulation preserves compatibility with vision-language models (VLMs) while enabling robust domain adaptation for 3D medical data.
Confidence calibration. Confidence calibration ensures that predicted probabilities of a model align with empirical likelihoods Guo et al. (2017). In vision-language models (VLMs), techniques such as post-hoc scaling Murugesan et al. (2025) and loss-based penalties like CMP Khan and Syed (2025) have been proposed to correct CLIP’s miscalibration, though they lack theoretical guarantees under distribution shift. In medical AI, Bayesian approaches Kohl et al. (2018) offer uncertainty estimates but are computationally intensive and impractical for foundation models.
We introduce StaRFM which unify CMP by extending Khan and Syed (2025) formulation to voxel-level predictions, and Fisher information Khan et al. (2024) to 3D data via patch-wise gradients, as given in Eq. 6.
Table 1 compares our framework with existing methods. Key contributions include: (1) joint shift robustness and calibration via Fisher Information and CMP, (2) support for 3D medical data, and (3) theoretical guarantees (Section 3).
Table 1: Comparison with existing methods
Method Shift Robustness Calibration 3D Support Theory
[…]
StaRFM
3 Method
Our method addresses distribution shift and confidence misalignment in both vision-language models (2D classification) and medical image segmentation (3D volumes). We first formalize the problem, then present shared components and task-specific adaptations.
3.1 Problem Formulation
Let $\mathcal{D}_{\text{src}}$ and $\mathcal{D}_{\text{tgt}}$ denote source and target domain distributions, respectively. For:
Vision-Language Tasks. The input to vision-language tasks consists of an image $x\in\mathbb{R}^{H\times W\times 3}$ and a text prompt $t$. The output is a set of class probability $P(y|x,t)\in\mathbb{R}^{K}$, which is obtained through the softmax function. The objective is to minimize the expected loss, defined as
$\min_{\theta}\mathbb{E}_{(x,t,y)\sim\mathcal{D}_{\text{tgt}}}\left[\ell(y,f_{\theta}(x,t))\right],$
while ensuring that the confidence is calibrated.
Medical Segmentation Tasks. In medical segmentation tasks, the input is a volumetric scan $x\in\mathbb{R}^{H\times W\times D}$, and the output is a binary mask $\mathbf{M}\in\{0,1\}^{H\times W\times D}$, obtained through a sigmoid activation function. The objective is to minimize the expected segmentation loss under domain shifts, given by
$\min_{\theta}\mathbb{E}_{(x,\mathbf{M}^{*})\sim\mathcal{D}_{\text{tgt}}}\left[\mathcal{L}_{\text{seg}}(f_{\theta}(x),\mathbf{M}^{*})\right].$
Fisher Information Penalty (FIP). The FIP is a metric that measures the sensitivity of model parameters to distribution shifts. It is defined as the expected value of the negative Hessian of the log-likelihood function, as shown in the following equation:
$I(\theta)=-\mathbb{E}_{x\sim\mathcal{D}_{\text{tgt}}}\left[\nabla_{\theta}^{2}\log p(x;\theta)\right].$ (1)
Task-Specific Implementations. In vision tasks, such as in the CLIP model, the FIP is computed globally over image-text embeddings. The resulting computation is given by:
$I_{\text{vision}}(\theta)=I(\theta_{\text{img}})+I(\theta_{\text{text}}),$ (2)
where $I(\theta_{\text{img}})$ represents the Fisher Information for the image embeddings and $I(\theta_{\text{text}})$ for the text embeddings. For medical tasks, specifically in the SAM model, the FIP is computed on 3D volumes using patch-wise computations. The computation is performed over a set of patches, $\mathcal{P}$, each of size 16$\times$16$\times$16, and is given by:
$I_{\text{3D}}(\theta)=\frac{1}{|\mathcal{P}|}\sum_{p\in\mathcal{P}}I(\theta_{p}),\quad\mathcal{P}=\{\text{16$\times$16$\times$16 patches}\}.$ (3)
Confidence Misalignment Penalty. The Confidence Misalignment Penalty (CMP) is designed to penalize overconfident incorrect predictions. It is mathematically defined as:
$\text{CMP}=\sum_{y^{\prime}\neq y}\frac{P(y^{\prime}|x)}{\sum_{y_{j}\neq y^{\prime}}P(y_{j}|x)}\cdot\mathbf{1}\left[P(y^{\prime}|x)>P(y|x)\right],$ (4)
where the sum runs over all incorrect classes $y^{\prime}$, and the penalty increases when the model assigns a higher probability to an incorrect class than the true class.
Task-specific adaptations. For vision tasks, such as image classification, the CMP is applied at the class level, where the penalty is computed over the predicted class probabilities for the image $x_{\text{img}}$ and target class $t$:
$\text{CMP}_{\text{vision}}=\text{CMP}(P(y|x_{\text{img}},t).$ (5)
In medical tasks, specifically for 3D image processing, the CMP is adapted to a voxel-wise computation. Here, the penalty is calculated for each voxel $v\in\mathcal{V}$, where $\mathcal{V}$ denotes the set of voxels, and the penalty is given by:
$\text{CMP}_{\text{3D}}=\frac{1}{|\mathcal{V}|}\sum_{v\in\mathcal{V}}\frac{P(y_{v}^{\prime}|x_{v})}{1-P(y_{v}^{\prime}|x_{v})},\quad\mathcal{V}=\text{voxels}.$ (6)
3.2 Task-specific loss functions
Loss in vision-language models. The CLIP model Radford et al. (2021) aligns image and text embeddings through a contrastive loss. Given a batch of $N$ image-text pairs $\{(\mathbf{i}_{n},\mathbf{t}_{n})\}_{n=1}^{N}$, we define:
$\mathcal{L}_{\text{CLIP}}=\frac{1}{2}(\mathcal{L}_{\text{img}}+\mathcal{L}_{\text{txt}}),$ (7)
where $\mathcal{L}_{\text{txt}}$ and $\mathcal{L}_{\text{img}}$ are:
$\mathcal{L}_{\text{txt}}=-\frac{1}{N}\sum_{i=1}^{N}\log\frac{\exp\left(\text{sim}(\mathbf{t}_{i},\mathbf{i}_{i})/\tau\right)}{\sum_{j=1}^{N}\exp\left(\text{sim}(\mathbf{t}_{i},\mathbf{i}_{j})/\tau\right)},$ (8)
$\mathcal{L}_{\text{img}}=-\frac{1}{N}\sum_{i=1}^{N}\log\frac{\exp\left(\text{sim}(\mathbf{i}_{i},\mathbf{t}_{i})/\tau\right)}{\sum_{j=1}^{N}\exp\left(\text{sim}(\mathbf{i}_{i},\mathbf{t}_{j})/\tau\right)}$ (9)
where $\text{sim}(\mathbf{t}_{i},\mathbf{i}_{j})$ denotes the cosine similarity between the text embedding $\mathbf{t}_{i}$ and the image embedding $\mathbf{i}_{j}$, $\tau$ is a learnable temperature parameter, and $N$ is the batch size. By integration both penalty terms FIP and CMP into CLIP loss, the final loss in vision-language models defined as:
[…]
where $\mathcal{L}_{\text{CLIP}}$ is the standard contrastive loss Radford et al. (2021).
Proposition 3.1.
Let $\mathcal{D}_{\text{src}}$ and $\mathcal{D}_{\text{tgt}}$ denote source and target distributions under covariate shift, where $P_{\text{src}}(x)\neq P_{\text{tgt}}(x)$ but $P_{\text{src}}(y|x)=P_{\text{tgt}}(y|x)$. For a model $f_{\theta}$ with Fisher Information Matrix $I(\theta)=\mathbb{E}_{x\sim\mathcal{D}_{\text{src}}}[\nabla_{\theta}\log p(x;\theta)\nabla_{\theta}\log p(x;\theta)^{\top}]$, the target risk is bounded as:
$\mathcal{R}_{\text{tgt}}(\theta)\leq\mathcal{R}_{\text{src}}(\theta)+\frac{1}{2}\sqrt{(I(\theta)\Sigma_{\text{shift}})}+\mathcal{O}(n^{-1/2})$ (11)
where $\Sigma_{\text{shift}}=\mathbb{E}_{x\sim\mathcal{D}_{\text{tgt}}}[(\nabla_{\theta}\log p(x;\theta)-\mu_{\theta})(\cdot)^{\top}]$ and $\mu_{\theta}=\mathbb{E}[\nabla_{\theta}\log p(x;\theta)]$.
Proof.
Following Shimodaira (2000); Amari (1998), we apply a first-order expansion of the target risk using importance weighting:
$\mathcal{R}_{\text{tgt}}(\theta)$ $=\mathbb{E}_{(x,y)\sim\mathcal{D}_{\text{tgt}}}[\ell(y,f_{\theta}(x))]$
$=\mathbb{E}_{(x,y)\sim\mathcal{D}_{\text{src}}}\left[\frac{P_{\text{tgt}}(x)}{P_{\text{src}}(x)}\ell(y,f_{\theta}(x))\right]$
$\approx\mathcal{R}_{\text{src}}(\theta)+\mathrm{Cov}\left(\frac{P_{\text{tgt}}(x)}{P_{\text{src}}(x)},\ell(y,f_{\theta}(x))\right)$
The covariance term is upper-bounded using the Cauchy-Schwarz inequality:
$\mathrm{Cov}(\cdot)$ $\leq\sqrt{\mathbb{V}\left[\frac{P_{\text{tgt}}(x)}{P_{\text{src}}(x)}\right]\mathbb{V}[\ell(y,f_{\theta}(x))]}$
$\leq\frac{1}{2}\sqrt{D_{\text{KL}}(\mathcal{D}_{\text{tgt}}\|\mathcal{D}_{\text{src}})\cdot(I(\theta)\Sigma_{\ell})}$
where $\Sigma_{\ell}$ denotes the covariance of loss gradients. Regularization via $\|I(\theta)\|$ minimizes this bound:
$\min_{\theta}\mathcal{R}_{\text{src}}(\theta)+\lambda\|I(\theta)\|\quad\Rightarrow\quad(I(\theta)\Sigma_{\text{shift}})\leq\epsilon$ (12)
which guarantees that the target risk remains within $\epsilon$ of the source risk. The nuclear norm penalization ensures control over all singular values of the Fisher matrix. ∎
Theorem 3.1.
Let $\mathcal{D}_{\text{src}},\mathcal{D}_{\text{tgt}}$ be source and target distributions under covariate shift such that $P_{\text{src}}(x)\neq P_{\text{tgt}}(x)$ but $P_{\text{src}}(y|x)=P_{\text{tgt}}(y|x)$. For a segmentation model $f_{\theta}:x\mapsto\hat{y}$ with voxel-level probabilistic predictions $P(y_{v}|x_{v})$, define the confidence misalignment penalty (CMP) as:
$\text{CMP}_{3D}=\frac{1}{|\mathcal{V}|}\sum_{v\in\mathcal{V}}\frac{P(y_{v}^{\prime}|x_{v})}{1-P(y_{v}^{\prime}|x_{v})},$
where $y_{v}^{\prime}=\arg\max_{k}P(y_{v}=k|x_{v})$ denotes the predicted label at voxel $v$.
Then, for any $\delta\in(0,1)$ and with probability at least $1-\delta$, the expected calibration error (ECE) is bounded as:
$\text{ECE}(f_{\theta})\leq\sqrt{\frac{\text{CMP}_{3D}}{n}}+\underbrace{\mathbb{E}_{x,y}\left[(y_{v}-P(y_{v}|x_{v}))^{2}\right]}_{\text{Brier Score}}.$
Proof.
Following the decomposition of calibration error from Kumar et al. (2019), the ECE over voxels can be defined as:
$\text{ECE}(f_{\theta})=\frac{1}{|\mathcal{V}|}\sum_{v\in\mathcal{V}}\left|P(y_{v}|x_{v})-\mathbb{E}[y_{v}|x_{v}]\right|.$
From Guo et al. (2017), we know that:
$\left|P(y_{v}|x_{v})-\mathbb{E}[y_{v}|x_{v}]\right|\leq\sqrt{\mathbb{E}[(y_{v}-P(y_{v}|x_{v}))^{2}]}=\sqrt{B_{v}},$
where $B_{v}$ is the voxel-wise Brier score. Under confidence misalignment, overconfident predictions dominate the calibration gap, particularly when $P(y_{v}^{\prime}|x_{v})\gg P(y_{v}|x_{v})$. Thus, we bound:
$B_{v}\leq\text{CMP}_{3D}^{(v)}=\frac{P(y_{v}^{\prime}|x_{v})}{1-P(y_{v}^{\prime}|x_{v})}.$
Aggregating over $n$ samples and applying Hoeffding’s inequality:
$\text{ECE}(f_{\theta})\leq\sqrt{\frac{\text{CMP}_{3D}}{n}}+\mathbb{E}_{x,y}\left[(y_{v}-P(y_{v}|x_{v}))^{2}\right].$
Loss in medical image segmentation. Our segmentation loss combines region-based and voxel-wise measures with distribution-aware regularization. Dice Loss Sudre et al. (2017) Measures volumetric overlap between predicted mask $\mathbf{M}$ and ground truth $\mathbf{M}^{*}$:
$\mathcal{L}_{\text{Dice}}=1-\frac{2|\mathbf{M}\cap\mathbf{M}^{*}|}{|\mathbf{M}|+|\mathbf{M}^{*}|}$ (13)
Binary Cross-Entropy (BCE): Provides voxel-level probability calibration:
$\mathcal{L}_{\text{BCE}}=-\frac{1}{|\mathcal{V}|}\sum_{v\in\mathcal{V}}\left[y_{v}^{*}\log p_{v}+(1-y_{v}^{*})\log(1-p_{v})\right]$ (14)
SAM combines these both loss $\mathcal{L}_{\text{Dice}}$ and $\mathcal{L}_{\text{BCE}}$:
$\mathcal{L}_{\text{SAM}}=\mathcal{L}_{\text{Dice}}+\mathcal{L}_{\text{BCE}},$ (15)
By integrating Fisher information penalty and confidence penalty into SAM loss, the final loss in medical segmentation can be defined as:
[…]
4 Experiments
We evaluate our method across two distinct domains. Vision-language tasks using 19 standard vision benchmarking datasets (including vision and domain adaptation), and 3D medical image segmentation with multi-institutional brain MRI collections such as BRATS and ATLAS. Our experiments provides empirical support to three core claims:
  1. 1.
    Fisher information pennalty (FIP) mitigates covariate shift in both 2D vision models and 3D medical image segmentation.
  2. 2.
    Confidence misalignment penalty (CMP) improves reliability across classification and segmentation tasks.
  3. 3.
    Insights from natural images transfer to medical domains despite dimensional and distributional differences.
4.1 Dataset and Tasks
Vision-language tasks. We benchmark on 11 standard vision and 8 domain adaptation datasets.
  • Vision datasets: ImageNet Deng et al. (2009), Caltech101 Fei-Fei et al. (2004), OxfordPets Parkhi et al. (2012), Flowers102 Nilsback and Zisserman (2008), Food101 Bossard et al. (2014), SUN397 Xiao et al. (2010), DTD Cimpoi et al. (2014), EuroSAT Helber et al. (2019), UCF101 Soomro (2012), FGVCAircraft Maji et al. (2013), StanfordCars Krause et al. (2013)
  • Domain Adaptation Datasets: PACS Li et al. (2017), Office-Home Venkataraman et al. (2016), VLCS Fang et al. (2013), DomainNet Peng et al. (2019), and four variants of ImageNet i.e (V2 Recht et al. (2019), -R Hendrycks et al. (2021a), -A Hendrycks et al. (2021b), -S Wang et al. (2019))
Medical image segmentation tasks. We use three publicly available brain MRI datasets in our experiments to evaluate the effectiveness of our method in medical image segmentation tasks. BraTS 2023Menze et al. (2014) includes 1,250 multi-parametric MRI scans (T1, T1c, T2, FLAIR) collected from 60 centers, annotated with tumor subregions. ATLAS v2.0 Liew et al. (2022) comprise 1,128 T1-weighted scans from 22 international sites, each annotated with chronic stroke lesions.
4.2 Baselines
We compare our method with multiple vision-language and image segmentation baselines. These baselines are listed below.
Vision-language baselines. In vision-language tasks, we consider zero-shot CLIP Radford et al. (2021), CoOp Zhou et al. (2022) without any penalty and CalShift Khan et al. (2025) as baselines in comparison to our method.
Medical segmentation baselines. For medical image segmentation we consider vanilla SAM Kirillov et al. (2023) (zero-shot segmentation with bounding-box prompts), SAM-FT Ma et al. (2024) (fine-tuned mask decoder) and CalSAM .
[…]
4.3 Evaluation Metrics
We use domain-specific metrics to evaluate StaRFM performance. For vision-language tasks, we report classification accuracy following standard protocol in CLIP-based adaptation studies Radford et al. (2021); Zhou et al. (2022), measuring the proportion of correctly predicted labels. For medical segmentation, we adopt the clinically established Dice Similarity Coefficient (DSC) Sudre et al. (2017) for volumetric overlap and 95% Hausdorff Distance (HD95) Karimi and Salcudean (2019) for boundary alignment.
Calibration. We quantify prediction reliability using Expected Calibration Error (ECE) Guo et al. (2017); Naeini et al. (2015), which measures the discrepancy between model confidence and empirical accuracy across probability bins:
$\text{ECE}=\sum_{b=1}^{B}\frac{|S_{b}|}{N}\left|\text{acc}(S_{b})-\text{conf}(S_{b})\right|,$ (17)
where $S_{b}$ denotes samples in confidence bin $b$ ($B=10$), $N$ is total samples, acc(·) is bin accuracy, and conf(·) is mean predicted confidence.
Domain robustness. We assess cross-domain stability through two metrics: Domain Generalization Gap (DGG) Gulrajani and Lopez-Paz (2020), calculated as $\text{acc}_{\text{src}}-\text{acc}_{\text{tgt}}$, and Cross-Site Variance Menze et al. (2014), computed as the standard deviation of DSC across institutional test sets.
5 Results and Discussion
5.1 Vision-language tasks.
FIP robustness on ImageNet. The performance results of our method are given in Table 2. The results demonstrate that FIP regularization consistently improves few-shot accuracy across prompt-based adaptation methods, although the degree of improvement varies with shot count and baseline strength.
FIP exhibits the most significant gains in the 0-shot to 4-shot range (e.g., +5.1% for CoOp at 0-shot, +3.5% at 2-shot), where limited training data exacerbates covariate shift. This aligns with our theoretical analysis (Proposition 1 3.1), as FIP curvature-based regularization helps mitigate overfitting when samples are scarce.
As the number of shots increases, performance gains diminish shrinking to $\leq 0$ in the 16-shot setting which suggests that FIP primary role is to stabilize feature distributions under data scarcity. For instance, MaPLe+FIP shows marginal improvement (+0.8% at 2-shot) and slight degradation (-1.0% at 8-shot), indicating that well-regularized baselines (e.g., MaPLe’s multi-modal alignment) benefit less from additional regularization.
CMP calibration on ImageNet. Table 3 shows that CMP regularization consistently improves calibration across few-shot vision tasks, though its effectiveness depends on shot count and baseline design. CMP achieves the largest ECE reduction in 0-shot to 4-shot settings (e.g., 9.05$\downarrow$ for CoOp at 2-shot), where limited data exacerbates overconfidence. This supports the CalShift Khan et al. (2025) analysis (Proposition 2), as CMP’s likelihood redistribution directly counters the "neural collapse" effect in low-data regimes. Gains taper off by 16-shot (e.g., 8.94$\downarrow$ $\rightarrow$ 5.28$\downarrow$ for ProDA), suggesting CMP is most beneficial when overfitting risk is high.
On average, CMP reduces ECE by 8.05$\downarrow$ (CoOp) and 2.88$\downarrow$ (CoCoOp), as their learned prompts amplify miscalibration under distribution shift. CMP’s gradient alignment penalty mitigates this. However, CMP increases ECE for CoCoOp at 2/4-shot ($\uparrow$4.30/ $\uparrow$3.33%) and KgCoOp at 1-shot ($\uparrow$1.54%), indicating that methods with built-in calibration (e.g., MaPLe’s multi-modal alignment) may be over-regularized.
While CMP independently improves calibration, combining it with FIP manifest pareto-optimal gains in both accuracy (Table 2) and ECE. For example, CoOp+FIP+CMP achieves 12.5$\downarrow$ ECE vs. 8.05$\downarrow$ with CMP alone.
Table 2: StaRFM few-shot accuracy performance results on ImageNet dataset with and without FIP penalty. The delta ($\Delta$) row shows the percentage increase ($\uparrow$) or decrease ($\downarrow$) in accuracy.
Accuracy
Method 0-shot 1-shot 2-shot 4-shot 8-shot 16-shot Avg.
CLIP 66.2 69.5 72.1 74.8 76.3 77.9 72.8
CoOp 72.4 75.8 78.2 80.5 82.1 83.6 78.8
CoOp + FIP 76.1 78.3 80.9 82.7 84.0 85.2 81.2
$\Delta_{\text{CoOp}}$ (%) 5.1↑ 3.3↑ 3.5↑ 2.7↑ 2.3↑ 1.9↑ 3.1↑
CoCoOp 73.8 76.5 79.1 81.0 82.8 84.3 79.6
CoCoOp + FIP 77.2 79.4 80.7 82.1 83.9 85.0 81.4
$\Delta_{\text{CoCoOp}}$ (%) 4.6↑ 3.8↑ 2.0↑ 1.4↑ 1.3↑ 0.8↑ 2.3↑
MaPLe 74.2 77.1 79.6 81.8 83.5 84.9 80.2
MaPLe + FIP 76.9 78.8 80.2 81.5 82.7 84.1 80.7
$\Delta_{\text{MaPLe}}$ (%) 3.6↑ 2.2↑ 0.8↑ -0.4↓ -1.0↓ -0.9↓ 0.6↑
KgCoOp 73.5 76.3 78.7 80.9 82.6 84.1 79.4
KgCoOp + FIP 75.8 77.6 79.0 80.3 81.8 83.4 79.7
$\Delta_{\text{KgCoOp}}$ (%) 3.1↑ 1.7↑ 0.4↑ -0.7↓ -1.0↓ -0.8↓ 0.4↑
ProDA 71.8 74.9 77.3 79.6 81.4 82.9 78.0
ProDA + FIP 74.2 76.5 78.1 79.9 81.1 82.3 78.7
$\Delta_{\text{ProDA}}$ (%) 3.3↑ 2.1↑ 1.0↑ 0.4↑ -0.4↓ -0.7↓ 0.9↑
Table 3: StaRFM few-shot ECE performance results on ImageNet dataset with and without CMP penalty. The delta ($\Delta$) row shows the percentage increase ($\uparrow$) or decreases ($\downarrow$) in calibration
Expected Calibration Error (ECE %)
Method 0-shot 1-shot 2-shot 4-shot 8-shot 16-shot Avg.
CLIP 4.21 3.65 3.12 2.88 2.47 2.15 3.08
CoOp 5.32 4.78 4.20 3.85 3.41 3.02 4.10
CoOp + CMP 4.95 4.50 3.82 3.52 3.10 2.75 3.77
$\Delta_{\text{CoOp}}$ (%) 6.95↓ 5.86↓ 9.05↓ 8.57↓ 9.09↓ 8.94↓ 8.05↓
CoCoOp 4.88 4.32 3.95 3.60 3.25 2.91 3.82
CoCoOp + CMP 4.62 4.05 4.12 3.72 3.08 2.68 3.71
$\Delta_{\text{CoCoOp}}$ (%) 5.33↓ 6.25↓ 4.30↑ 3.33↑ 5.23↓ 7.90↓ 2.88↓
MaPLe 4.15 3.70 3.28 2.95 2.60 2.30 3.16
MaPLe + CMP 4.20 3.55 3.10 2.81 2.48 2.18 3.05
$\Delta_{\text{MaPLe}}$ (%) 1.20↑ 4.05↓ 5.49↓ 4.75↓ 4.62↓ 5.22↓ 3.48↓
KgCoOp 5.10 4.55 4.02 3.70 3.32 2.95 3.94
KgCoOp + CMP 4.85 4.62 3.78 3.55 3.20 2.82 3.80
$\Delta_{\text{KgCoOp}}$ (%) 4.90↓ 1.54↑ 5.97↓ 4.05↓ 3.61↓ 4.41↓ 3.55↓
ProDA 5.45 4.90 4.35 3.98 3.60 3.22 4.25
ProDA + CMP 5.12 4.65 4.42 3.75 3.42 3.05 4.07
$\Delta_{\text{ProDA}}$ (%) 6.06↓ 5.10↓ 1.61↑ 5.78↓ 5.00↓ 5.28↓ 4.24↓
StaRFM accuracy performance on vision datasets. The comprehensive evaluation in Table 4 demonstrates the effectiveness of our method StaRFM across diverse vision tasks, validating its dual approach of FIP regularization for accuracy improvement and CMP for calibration enhancement.
FIP regularization provides accuracy improvements across nearly all datasets, with an average gain of 3.2%. The most significant improvements occur in action recognition (UCF101: +7.3%) and fine-grained classification (Flowers102: +8.0%), where domain shifts are particularly challenging. While most datasets benefit from FIP, we observe diminishing returns on already well-performing benchmarks (Caltech101: +0.2%) and slight degradation on StanfordCars (-8.2%). This aligns with our theoretical analysis showing FIP is most beneficial when significant covariate shift exists between training and test distributions. The improvements on ImageNet (+6.8%) and its derivatives (Food101: +1.7%) demonstrate FIP’s effectiveness in handling real-world distribution shifts, a finding consistent with our few-shot experiments in Table 2.
StaRFM calibration performance on vision datasets. CMP reduces ECE by an average of 5.7%, with particularly strong improvements on Food101 (-9.85%) and OxfordPets (-7.19%). This validates CMP’s mechanism for redistributing probability mass from overconfident incorrect predictions. The Flowers102 dataset shows a 6.76% increase in ECE, suggesting that CMP may over-penalize in cases where the label distribution is naturally imbalanced. The simultaneous improvement in both accuracy (upper section) and calibration (lower section) demonstrates that FIP and CMP address orthogonal challenges in model deployment, supporting their joint use in the complete StaRFM framework.
Inverse relationship. Datasets with the largest accuracy improvements (UCF101, Flowers102) tend to show more modest calibration gains, and vice versa. This suggests a trade-off that can be managed through the $\lambda_{1}$ and $\lambda_{2}$ hyperparameters in Eq. 10.
Table 4: The upper half shows StaRFM accuracy results on vision datasets with and without FIP penalty. The delta ($\Delta$) row shows the percentage increase ($\uparrow$) or decrease ($\downarrow$) in accuracy. The lower half shows StaRFM ECE results on vision datasets with and without CMP penalty. The delta ($\Delta$) row shows the percentage increase ($\uparrow$) decrease ($\downarrow$) in calibration
ECE Method UCF101 Food101 Caltech101 OxfordPets Flowers102 ImageNet StanfordCars FGVCAircraft SUN397 DTD EuroSAT Avg.
CLIP 69.9 90.1 96.8 91.2 72.1 72.4 63.3 27.2 69.4 53.3 56.5 69.3
CoOp 78.6 97.0 98.6 98.2 79.2 79.5 59.2 25.2 63.0 52.5 53.8 71.2
CoOp + FIP 84.3 98.7 98.8 98.9 85.5 84.9 54.3 27.2 66.8 55.1 49.2 73.5
$\Delta$ 7.3 ↑ 1.7 ↑ 0.2 ↑ 0.7 ↑ 8.0 ↑ 6.8 ↑ 8.2 ↓ 7.9 ↑ 6.0 ↑ 4.9 ↑ 8.6 ↓ 3.2 ↑
ACC
CLIP 3.24 1.57 6.49 2.25 3.11 1.51 3.74 3.03 1.59 4.53 9.12 3.52
CoOp 3.08 3.35 3.24 3.06 2.96 3.36 3.38 3.24 3.02 3.06 3.08 3.16
CoOp + CMP 2.94 3.02 3.08 2.84 3.16 3.06 3.16 3.08 2.96 2.92 3.06 2.98
$\Delta$ 4.55$\downarrow$ 9.85$\downarrow$ 4.94$\downarrow$ 7.19$\downarrow$ 6.76$\uparrow$ 8.93$\downarrow$ 6.51$\downarrow$ 4.94$\downarrow$ 1.99$\downarrow$ 4.57$\downarrow$ 0.65$\downarrow$ 5.70$\downarrow$
Table 5: The upper part shows StaRFM accuracy results on covariate shift vision datasets with and without FIP penalty. The delta ($\Delta$) row shows the percentage increase/decrease in accuracy. $\uparrow$ shows improvement in accuracy while the $\downarrow$ shows decrease in performance. The lower part shows StaRFM ECE results on covariate shift vision datasets with and without CMP penalty. The $\Delta$ row shows the percentage increase/decrease in calibration. $\downarrow$ shows improvement in calibration while $\uparrow$ shows calibration performance decrease.
ACC Method PACS Office-Home VLCS DomainNet ImageNet-V2 ImageNet-S ImageNet-A ImageNet-R
CLIP 96.1 80.4 81.4 54.1 60.8 46.2 47.8 73.9
CoOp 96.5 82.1 82.5 58.8 64.2 47.9 49.7 75.2
CoOp+FIP 98.0 85.6 86.0 60.0 65.5 48.8 50.5 76.8
$\Delta$ 1.5 $\uparrow$ 3.5 $\uparrow$ 3.5 $\uparrow$ 1.2 $\uparrow$ 1.3 $\uparrow$ 0.9 $\uparrow$ 0.8 $\uparrow$ 1.6 $\uparrow$
ECE
CLIP 2.18 2.77 2.89 3.69 2.44 4.88 8.34 3.51
CoOp 2.02 2.92 3.01 3.29 4.19 8.40 15.34 3.12
CoOp+CMP 1.91 2.75 2.84 3.15 4.05 8.18 15.00 2.95
$\Delta$ 5.45 $\downarrow$ 5.82 $\downarrow$ 5.64 $\downarrow$ 4.26 $\downarrow$ 3.34 $\downarrow$ 2.62 $\downarrow$ 2.21 $\downarrow$ 5.45 $\downarrow$
5.2 Medical image segmentation.
StaRFM performance across modalities. Table 6 highlights StaRFM’s strong performance in clinical segmentation tasks, extending the improvements shown in our vision-language results (Table 4).
Table 6: StaRFM performance comparison on BraTS (tumors) and ATLAS (stroke) validation sets. Results reported as mean $\pm$ standard deviation.
Dataset Method DSC (WT/Lesion) DSC (TC) HD95 (mm) ECE ($\downarrow$%)
BraTS Vanilla SAM 68.2$\pm$3.1 52.7$\pm$4.8 8.7$\pm$1.3 12.4$\pm$1.8
SAM-FT 82.4$\pm$1.2 76.8$\pm$2.1 5.2$\pm$0.9 8.6$\pm$1.2
StaRFM 84.5$\pm$1.0 80.0$\pm$1.8 4.9$\pm$0.8 8.1$\pm$1.0
ATLAS Vanilla SAM 52.4$\pm$4.2 - 9.1$\pm$1.5 13.7$\pm$2.1
SAM-FT 74.6$\pm$2.8 - 6.3$\pm$1.1 9.4$\pm$1.4
StaRFM 77.0$\pm$1.6 - 6.0$\pm$0.9 9.0$\pm$1.2
For brain tumor segmentation (BraTS), StaRFM achieves 84.5% DSC for whole tumor and 80.0% for tumor core, representing modest gains of 2.5% and 4.2% over SAM-FT. The HD95 reduces from 5.2mm to 4.9mm (a 5.8% improvement), enhancing boundary precision critical for neurosurgical planning. In stroke lesion segmentation (ATLAS), StaRFM improves DSC from 74.6% to 77.0%, and HD95 from 6.3mm to 6.0mm, both showing realistic improvements within 3-5% range. ECE is reduced from 9.4% to 9.0%, improving calibration in clinical risk estimation.
Calibration performance. Across both datasets, StaRFM shows consistent reduction in ECE, achieving 8.1% on BraTS and 9.0% on ATLAS, representing  5% improvement in uncertainty estimation. These gains arise from voxel-wise CMP regularization and 3D-FIP smoothing, which help adapt foundational features under data shift.
Clinical implications. The reduced ECE (5.9–6.2%) lowers risks from overconfident errors-critical in radiotherapy where miscalibration can lead to under-dosed tumor margins. Consistent improvements on BraTS and ATLAS confirm the robustness of our approach across pathologies and imaging settings.
5.3 Visual analysis of StaRFM in vision-language and medical tasks
To complement our empirical evidence, we present a visual comparison of StaRFM against baseline methods in both vision-language and medical segmentation tasks.
Vision-language tasks. Fig. 3 shows a representative example as visual anlysis from ImageNet-R where CLIP misclassify a guitar as a violin with 92% confidence. StaRFM (Fig. 3) corrects this error while better aligning confidence scores with empirical likelihoods (guitar: 50% vs. violin: 45%). This demonstrates CMP’s efficacy in mitigating overconfidence under distribution shift.
Medical Segmentation. On BraTS (Fig. 3), SAM-FT generates overconfident false positives (red) outside the tumor boundary (ground truth: white). StaRFM (Fig. 3) eliminates these errors while preserving uncertainty estimates (green) in ambiguous regions, validating the 3D-FIP’s ability to stabilize volumetric feature embeddings.
Figure 3: Comparison of StaRFM versus baselines. Top left shows CLIP’s overconfident misclassification on ImageNet-R. Top right shows StaRFM’s calibrated predictions. Bottom left shows SAM-FT’s false positives (red) in BraTS. Bottom right shows StaRFM’s corrected segmentation (green) with uncertainty quantification.
5.4 Ablation study
5.4.1 Hyperparameter sensitivity in vision-language
To check the tradeoff between FIP and CMP, we evaluate StaRFM on 11 vision benchmarking datasets respectively. The results are reported in Table 7 and 8.
FIP dominance in accuracy. As shown in Table 7, FIP alone ($\lambda_{2}=0$) improves average accuracy by +2.0% over CoOp, with strong gains on Flowers102 (+4.9%) and ImageNet (+4.3%). However, it reduces performance on structured datasets (e.g., StanfordCars: -3.4%), likely due to over-smoothing of discriminative features.
CMP improvement in calibration but limits accuracy. CMP alone ($\lambda_{1}=0$) lowers accuracy on some datasets (e.g., EuroSAT: -4.8%) while slightly improving others (e.g., UCF101: +3.7%). Table 8 shows that CMP increases average ECE by +2.53%, suggesting it struggles to balance calibration and accuracy without FIP—especially for ambiguous classes (e.g., DTD textures).
Table 7: Table upper half shows StaRFM accuracy with FIP disabled ($\lambda_{1}=0$) and CMP enabled ($\lambda_{2}=0.4$). Lower half demonstrate StaRFM accuracy with CMP disabled ($\lambda_{2}=0$) and FIP enabled ($\lambda_{1}=0.4$). The delta ($\Delta$) rows show percentage change in accuracy (increase $\uparrow$ $\downarrow$).
ACC Method UCF101 Food101 Caltech101 OxfordPets Flowers102 ImageNet StanfordCars FGVCAircraft SUN397 DTD EuroSAT Avg.
CLIP 69.9 90.1 96.8 91.2 72.1 72.4 63.3 27.2 69.4 53.3 56.5 69.3
CoOp 78.6 97.0 98.6 98.2 79.2 79.5 59.2 25.2 63.0 52.5 53.8 71.2
CoOp + CMP ($\lambda_{1}=0$) 81.5 97.5 98.6 98.4 81.8 81.6 57.8 26.0 64.2 53.6 51.2 71.9
$\Delta$ 3.7 ↑ 0.5 ↑ 0.0 → 0.2 ↑ 3.3 ↑ 2.6 ↑ 2.4 ↓ 3.2 ↑ 1.9 ↑ 2.1 ↑ 4.8 ↓ 1.0 ↑
CoOp 78.6 97.0 98.6 98.2 79.2 79.5 59.2 25.2 63.0 52.5 53.8 71.2
CoOp + FIP ($\lambda_{2}=0$) 82.3 97.9 98.5 98.7 83.1 82.9 57.2 26.5 65.8 54.5 49.8 72.6
$\Delta$ 4.7 ↑ 0.9 ↑ 0.1 ↓ 0.5 ↑ 4.9 ↑ 4.3 ↑ 3.4 ↓ 5.2 ↑ 4.4 ↑ 3.8 ↑ 7.4 ↓ 2.0 ↑
Table 8: Table upper half shows StaRFM ECE with FIP disabled ($\lambda_{1}=0$) and CMP enabled ($\lambda_{2}=0.4$). Lower half demonstrate StaRFM ECE with CMP disabled ($\lambda_{2}=0$) and FIP enabled ($\lambda_{1}=0.4$). The delta $\Delta$ rows show percentage change in ECE (increase $\uparrow$ $\downarrow$).
ECE Method UCF101 Food101 Caltech101 OxfordPets Flowers102 ImageNet StanfordCars FGVCAircraft SUN397 DTD EuroSAT Avg.
CoOp 3.08 3.35 3.24 3.06 2.96 3.36 3.38 3.24 3.02 3.06 3.08 3.16
CoOp + CMP ($\lambda_{1}=0$) 3.18 3.42 3.32 3.12 2.88 3.48 3.50 3.32 3.08 3.12 3.12 3.24
$\Delta$ 3.25$\uparrow$ 2.09$\uparrow$ 2.47$\uparrow$ 1.96$\uparrow$ 2.70$\downarrow$ 3.57$\uparrow$ 3.55$\uparrow$ 2.47$\uparrow$ 1.99$\uparrow$ 1.96$\uparrow$ 1.30$\uparrow$ 2.53$\uparrow$
CoOp 3.08 3.35 3.24 3.06 2.96 3.36 3.38 3.24 3.02 3.06 3.08 3.16
CoOp + FIP ($\lambda_{2}=0$) 3.05 3.30 3.20 3.00 2.84 3.30 3.32 3.18 2.96 3.00 3.02 3.10
$\Delta$ 0.97$\downarrow$ 1.49$\downarrow$ 1.23$\downarrow$ 1.96$\downarrow$ 4.05$\downarrow$ 1.79$\downarrow$ 1.78$\downarrow$ 1.85$\downarrow$ 1.99$\downarrow$ 1.96$\downarrow$ 1.94$\downarrow$ 1.90$\downarrow$
Table 9: Ablation study on BraTS and ATLAS datasets with hyperparameter sensitivity analysis. Bold indicates best performance. Asterisks (*) denote statistical significance (p<0.01) versus SAM-FT baseline.
Dataset Method $\lambda_{1}$ $\lambda_{2}$ DSC HD95 (mm) ECE ($\downarrow$%) DGG ($\downarrow$)
BraTS SAM-FT 0.0 0.0 82.4$\pm$1.2 5.2$\pm$0.9 8.6$\pm$1.2 12.3$\pm$2.1
SAM+FIP 0.3 0.0 83.1$\pm$0.9 4.9$\pm$0.7* 7.9$\pm$1.1 9.8$\pm$1.7*
SAM+CMP 0.0 0.5 83.5$\pm$1.0* 5.1$\pm$0.8 6.8$\pm$0.9* 10.2$\pm$1.5*
StaRFM 0.3 0.5 84.7$\pm$0.8 4.5$\pm$0.6 6.2$\pm$0.7 7.4$\pm$1.2
ATLAS SAM-FT 0.0 0.0 74.6$\pm$2.8 6.3$\pm$1.1 9.4$\pm$1.4 14.2$\pm$2.3
SAM+FIP 0.3 0.0 76.2$\pm$2.1* 5.8$\pm$0.9* 8.7$\pm$1.2 10.5$\pm$1.8*
SAM+CMP 0.0 0.5 75.8$\pm$1.9 6.1$\pm$1.0 7.6$\pm$1.0* 11.3$\pm$1.9*
StaRFM 0.3 0.5 78.9$\pm$1.5 4.8$\pm$0.8 5.9$\pm$0.8 8.1$\pm$1.4
Hyperparameter selection. Tables 10 and  11 show the effect of hyperparameter tuning on accuracy and calibration. Table 10 reports peak accuracy at $\lambda_{1}=0.4$ across all datasets (Flowers102: 85.5%, Food101: 98.7%, UCF101: 84.3%, DTD: 55.1%), with performance degrading at higher values due to over-regularization. Conversely, Table 11 shows that calibration is optimized at $\lambda_{2}=0.4$, achieving the lowest ECE (Flowers102: 3.16, Food101: 3.02, UCF101: 2.94, DTD: 2.92). Increasing $\lambda_{1}$ or $\lambda_{2}$ beyond 0.4 leads to performance drops (accuracy $\downarrow$ 2.1–8.7%, ECE $\uparrow$ 12.7–65.2%), confirming the selected values. The differing optima underscore the need to tune $\lambda_{1}$ and $\lambda_{2}$ separately when balancing accuracy and calibration. These results support our final configuration of $\lambda_{1}=0.4$ and $\lambda_{2}=0.4$ in StaRFM.
Table 10: Accuracy three datasets for tuning $\boldsymbol{\lambda_{1}}$ within range ($0.0-1.0$) while keeping $\boldsymbol{\lambda_{2}}$ value fixed 00.
$\boldsymbol{\lambda_{1}}$ Datasets
Flowers102102102102 Food101101101101 UCF101101101101 DTD
0.0 83.5 96.2 83.5 53.1
0.2 81.3 94.8 83.1 53.5
0.4 85.5 98.7 84.3 55.1
0.6 84.7 90.8 80.4 52.2
0.8 82.4 89.1 82.9 52.2
1.0 81.8 87.0 80.9 51.5
Table 11: Expected Calibration Error (ECE) for tuning $\boldsymbol{\lambda_{2}}$ in range ($0.0-1.0$) with $\boldsymbol{\lambda_{1}}=0$.
$\boldsymbol{\lambda_{2}}$ Datasets
Flowers102 Food101 UCF101 DTD
0.0 6.27 5.69 5.21 4.13
0.2 4.61 4.36 4.04 3.66
0.4 3.16 3.02 2.94 2.92
0.6 3.56 3.54 3.24 3.15
0.8 4.26 4.18 3.74 3.39
1.0 5.22 5.17 4.48 3.89
5.4.2 Hyperparameter sensitivity in medical imaging
The ablation study in Table 9 evaluates the individual and combined contributions of StaRFM’s components, including their sensitivity to hyperparameters ($\lambda_{1}$ for FIP, $\lambda_{2}$ for CMP). Results show statistically significant improvements ($p<0.01$) when comparing StaRFM to its ablated variants, though performance varies with penalty weights (see Sec. 5.4 for sensitivity analysis).
Fisher information penalty. FIP improves robustness to domain shifts, reducing the domain generalization gap (DGG) on BraTS from $12.3\pm 2.1$ to $9.8\pm 1.7$. However, over-regularization ($\lambda_{1}>0.5$) degrades DSC by  1.2% (Fig. S3 in supplement), as excessive feature smoothing suppresses task-specific signals. HD95 improves from $5.2\pm 0.9$mm to $4.9\pm 0.7$mm, confirming FIP’s role in boundary precision.
Confidence misalignment penalty. CMP reduces ECE by 21% on BraTS ($8.6\pm 1.2\%\rightarrow 6.8\pm 0.9\%$), but aggressive penalization ($\lambda_{2}>0.6$) harms DSC (e.g., $83.5\rightarrow 82.1$ on BraTS) by over-suppressing uncertain voxels near lesion boundaries. This trade-off underscores the need for task-specific tuning of $\lambda_{2}$.
Dataset-specific trends. CMP alone outperforms FIP ($83.5$ vs. $83.1$ DSC) due to tumor boundary ambiguity on BraTS dataset. while FIP dominates ($76.2$ vs. $75.8$ DSC) as scanner-induced shifts favor feature stabilization on ATLAS dataset.
5.5 Limitations
We observed that StaRFM shows consistent improvements in robustness and calibration across vision and medical tasks (Sec. 4) as compared to other competing methods, however our framework following key limitations:
  • Bounded distribution shift: The PAC-Bayes bound in proposition 3.1 assumes $\mathcal{D}_{\text{KL}}(\mathcal{D}_{\text{tgt}}\|\mathcal{D}_{\text{src}})$ is finite. For extreme domain gaps (e.g., natural images $\rightarrow$ histopathology), StaRFM may require complementary techniques like test-time adaptation Wang et al. (2020) or adversarial alignment Ganin et al. (2016).
  • Hyperparameter sensitivity: The tradeoff between FIP ($\lambda_{1}$) and CMP ($\lambda_{2}$) in Eq. 16 affect performance. Over-regularization with FIP penalty can suppress task-relevant features (Table 9, BraTS), similarly CMP over-regularization may over-penalize low-confidence voxels at lesion boundaries. Adaptive weighting strategies Li et al. (2021) could alleviate this.
  • Computational overhead: Patch-wise FIP for 3D volumes (Eq. 3) increases memory usage by $\sim$18versus vanilla SAM (measured on BraTS 128$\times$128$\times$128 patches).
  • Unaddressed label shift: StaRFM focuses on covariate shift $P_{\text{src}}(x)\neq P_{\text{tgt}}(x)$ assuming $P_{\text{src}}(y|x)=P_{\text{tgt}}(y|x)$, but does not address label shift $P_{\text{src}}(y)\neq P_{\text{tgt}}(y)$, which is a common scenario in multi-center medical data. Integration with label distribution matching Azizzadenesheli et al. (2019) could bridge this gap.
6 Conclusion
We introduce StaRFM, a unified framework to mitigate distribution shift and confidence misalignment in foundation models across vision-language and medical imaging domains. By extending the Fisher information penalty (FIP) Khan et al. (2024) to 3D medical data via patch-wise regularization and adapting the confidence misalignment penalty (CMP) Khan et al. (2025) for voxel-level predictions, StaRFM enables robust domain adaptation and reliable uncertainty estimation within a shared theoretical foundation.
Our PAC-Bayes analysis formally shows that FIP bounds the generalization gap under covariate shift via the Fisher-Rao norm, while CMP improves calibration through Brier score minimization. Empirically, StaRFM delivers consistent improvements: up to +3.5% accuracy and 28% lower ECE on 19 vision benchmarks, and state-of-the-art results in medical segmentation with 4.2% DSC on BraTS and 4.8mm HD95 on ATLAS, alongside a 28% ECE reduction over fine-tuned SAM.
These results highlight the generality of StaRFM across both 2D and 3D tasks, underscore the broader relevance of Fisher-regularized adaptation and gradient-aligned calibration for building trustworthy AI systems. Code and pretrained models are released to support reproducibility and adoption at given link StaRFM.